OpenStudy (anonymous):

x2 + 5x + 22 = 0 I got :-2.5 am i right??

OpenStudy (anonymous):

NO the answer is x=-4.4

OpenStudy (anonymous):

yes U are right

OpenStudy (anonymous):

-2.5 is right??

OpenStudy (anonymous):

em sure it is right

OpenStudy (anonymous):

it's all wrong. \[D=b^{2}−4∗a∗c=25−4∗1∗22=−63\] D<0 , so there will be complex roots (solutions), given by the formulas: \[x_{1,2}=(−b±i∗\sqrt{-D})/(2∗a) = (-5 \pm i*\sqrt{63})/2\] , or by common formula with no calculating of \[\sqrt{-63}\] , i.e. \[x_{1,2}=(−b±\sqrt{-D})/(2∗a)=(−5±\sqrt{−63})/2\]